# Dual Test-time Training for Out-of-distribution Recommender System

### v1

@@ Abstract
Deep learning has been widely applied in recommender systems, which has achieved revolutionary progress recently. However, most existing learning-based methods assume that the user and item distributions remain unchanged between the training phase and the test phase. However, the distribution of user and item features can naturally shift in real-world scenarios, potentially resulting in a substantial decrease in recommendation performance. This phenomenon can be formulated as an Out-Of-Distribution (OOD) recommendation problem. To address this challenge, we propose a novel Dual Test-Time-Training framework for OOD Recommendation, termed DT3OR. In DT3OR, we incorporate a model adaptation mechanism during the test-time phase to carefully update the recommendation model, allowing the model to specially adapt to the shifting user and item features. To be specific, we propose a self-distillation task and a contrastive task to assist the model learning both the user’s invariant interest preferences and the variant user/item characteristics during the test-time phase, thus facilitating a smooth adaptation to the shifting features. Furthermore, we provide theoretical analysis to support the rationale behind our dual test-time training framework. To the best of our knowledge, this paper is the first work to address OOD recommendation via a test-time-training strategy. We conduct experiments on three datasets with various backbones. Comprehensive experimental results have demonstrated the effectiveness of DT3OR compared to other state-of-the-art baselines.
@@ I Introduction
Recommender systems play a crucial role in alleviating information overload on social media platforms by providing personalized information filtering. In recent years, a plethora of recommendation algorithms have been proposed, including collaborative filtering [1, 2, 3, 4, 5, 6], graph-based recommendation [7, 8, 9], cross-domain recommendation [10] and etc.
Although promising performance has been achieved, most existing recommendation algorithms assume that the user and item distributions remain unchanged between the training phase and the test phase, adhering to the assumption of independent and identically distributed (I.I.D.) variables. However, these distributions often undergo shifts in many real-world scenarios, violating this assumption. The shifts primarily manifest in two aspects: user feature shift and item feature shift. User feature shift specifically refers to changes in a user’s preferences over time as his or her characteristics evolve. For example, as a boy grows older, his interests may shift from comic books to sports, such as basketball. On the other hand, item feature shift pertains to changes in preference relationships resulting from the replacement or updating of specific types of items. For instance, when items are renewed or updated, they may have new characteristics that lead to different preferences and referral relationships among users. We demonstrate the user/item feature shift in Fig. 1. These shifts significant impact on recommendation performance [11, 12]. Consequently, it is crucial to consider out-of-distribution (OOD) recommendation to address this issue.
@@
The out-of-distribution (OOD) problem has been extensively studied in various fields, including image classification [13] and detection [14]. However, it has received limited attention in the context of recommendation systems. To address this issue, disentangled recommendation approaches [15, 16] aim to learn factorized representations of user preferences, which can enhance the model robustness against distribution shifts. Causal-based methods [11, 12] use causal learning to solve OOD problems. However, these methods often require interventions during the training process, which makes them less practical when only a pre-trained model is available. Another approach to address OOD recommendation is model retraining [8], which involves adapting the model to the OOD environment. However, this approach often requires a substantial computational overhead. [11]. In contrast, our study aims to investigate practical approaches that can readily improve the generalization and robustness for a wide range of pre-trained models in different scenarios. Test-time training (TTT), a widely adopted strategy for solving OOD problems in various domains [17], serves as the inspiration for our proposed solution. The core concept of test-time training is to utilize auxiliary unsupervised or self-supervised learning (SSL) tasks to provide insights into the underlying characteristics of a test sample. This information is then used to fine-tune or calibrate the trained model specifically for that test sample. Since each sample has distinct properties, test-time training has significant potential to enhance the generalization capabilities of trained models.
@@
  • To the best of our knowledge, this paper is the first work to deal OOD recommendation problem with test-time-training, which allows the model to adapt for the shifting user/item features.
  • To learn user-invariant preference relationships and enable the model to adapt to shifting user and item features, we propose a dual test-time training strategy that incorporates two self-supervised learning (SSL) tasks: the self-distillation task and the contrastive task.
  • We provide a theoretical analysis to demonstrate the rationale behind our designed dual test-time-training strategy.
  • Comprehensive experimental evidence across three datasets has shown that DT3OR surpasses other state-of-the-art recommendation algorithms in OOD scenario. In addition, the efficiency of the proposed modules is confirmed through ablation studies.
@@ II Problem Definition
III Methodology
@@
We design a novel framework termed DT3OR for recommendation under an OOD scenario in this section. The framework of DT3OR is shown in Fig. 2. To be specific, we first describe the overall framework of DT3OR. Then, we delve into the designed self-supervised learning tasks, which include the self-distillation task and the contrastive task. Finally, we give the definition of loss function in DT3OR.
The primary obstacle in optimizing the DT3OR problem lies in constructing a surrogate loss that can guide the learning process, particularly considering the lack of ground-truth interactions for the test data. To tackle this challenge, we present a dual self-supervised learning task as delineated below.
III-A Self-supervised Learning Tasks Designing
@@
Our proposed DT3OR is designed to enhance model generalization and robustness to handle the OOD recommendations by learning from the test data. Ideally, if we had the ground-truth labels for the test data, we could refine the model to minimize the cross-entropy loss on the test data. However, due to the absence of such label information at the test-time, we are motivated to explore suitable surrogate loss functions to steer the model without the presence of labeled data. In the unsupervised scenario, self-supervised learning (SSL) techniques [18, 19, 20] have gained significant attention, laying the groundwork for self-supervision during test-time training. The effectiveness of test-time training heavily dependents on the creation of a relevant and informative SSL task [17]. However, SSL tasks commonly used in image-related tasks, such as rotation invariance, may not be directly applicable to recommendation data due to its distinct characteristics. In following, we design two SSL tasks that are suitable for recommendation tasks.
III-A1 Self-distillation Task Designing
@@
In this subsection, we propose a test-time self-distillation task to make better leverage of the test data. To be specific, we first extract the embeddings of user/item features and historical interactions by the pretrain user feature network $f(\cdot)$ and item feature network $g(\cdot)$:
@@
where $\textbf{X}_{u}$ and $\textbf{X}_{i}$ denote the user features and item features, respectively. D is the historical interactions. $\textbf{E}_{u}$ and $\textbf{E}_{i}$ capture the preferences of users towards items. In a recommendation system, historical interactions and user features are valuable indicators of user preferences [21]. User’s purchase behaviors could reflect their preference for a particular type of item. Similarly, the items interacted with by users can indicate their preferences.
@@ E $=\gamma(\textbf{E}_{u}\oplus\textbf{E}_{i}),$ (3)
where $\oplus$ denotes the the concatenation operation, $\gamma(\cdot)$ denotes a multi-layer perceptron (MLP). We implement K-means algorithm [22] on E. The clustering results distribution could be obtained by:
@@
where $\varphi(\cdot)$ denotes the K-means algorithm [22]. The variable $q$ can be interpreted as the probability of assigning a sample to a specific cluster, indicating a soft assignment. We can consider the matrix $\textbf{Q}=[q]$ as the distribution of assignments for all samples.
@@
III-A2 Contrastive Task Designing
Contrastive learning has surfaced as a significant self-supervised framework, demonstrating promising performance across various tasks. In this subsection, we design a contrastive task specifically for test-time training. Data augmentation is pivotal in contrastive learning. However, inappropriate data augmentation can lead to semantic drift, thereby degrading the quality of positive samples [23]. Additionally, existing strategies for constructing negative samples often consider all non-positive samples as negative, which can inadvertently introduce false-negative samples [24].
In this subsection, we design a contrastive task with more discriminative positive samples and reliable negative samples. Based on the clustering results obtained by Eq.(4), we design a contrastive task in this section. To be concrete, the designed contrastive task contains two steps.
Firstly, we construct the positive sample pairs. Based on the clustering distribution $q$, we identify the high confidence sample indexes as:
@@ III-A2 Contrastive Task Designing
where $\textbf{C}_{i}$ represents different preference centers. We consider different high-confidence centers as negative samples, as they correspond to different interest centers. In this way, we could reduce the possibility of false-negative samples.
@@ III-B Objective function
The first term in Eq.(14) promotes the distribution of user interest centers closer together, enhancing the uniformity of user interest representations in the latent space. The second term captures the correlations among users with similar preferences. Here, we use ${B}$ to represent the batch size. Furthermore, the embeddings’ dimension is set to $d$, and $N$ as the number of high-confidence samples. Consequently, we could obtain time complexity of the self-distillation loss and contrastive loss is $\mathcal{O}(Bd)$ and $\mathcal{O}(N^{2}d)$, respectively. Therefore, the total time complexity of the loss is $\mathcal{O}(Bd+N^{2}d)$. The space complexity of our loss is $\mathcal{O}(B+N^{2})$. The comprehensive learning process of DT3OR is outlined in Algorithm 1.
IV Theoretical analysis
@@
With the inner product of embeddings $g=f(u,i)$ and based on Eq (15), we could have the following formula:
@@
where $\hat{y}_{i}$ is the prediction of the recommendation model. Then, we calculate the Hessian Matrix of Eq (17):
@@
Based on Eq.(IV), for any ${\bf c}\in\mathbb{R}^{N}$, we could have:
@@
Therefore, we have successfully demonstrated that the Hessian Matrix $\mathcal{\bf M}$ of Eq (17) is positive semi-definite (PSD). This proof is equivalent to establishing the convexity of $\mathcal{L}(y;{\bf g})$ with respect to $\bf g$.
@@
Remark. By applying Theorem1, we can assert that the objective loss function of the recommendation model possesses the properties of differentiability, convexity, and $\beta$-smoothness. Furthermore, in combination with TheoremIV, we can conclude that test-time training could enhance the performance of the recommender system.
@@
Table I: Statistics of the datasets used in our method. Besides, ”int” represents interactions.
Dataset Synthetic Meituan Yelp
#User 1,000 2,145 7,975
#Item 1,000 7,189 74,722
#OOD Int 112,371 6,944 99,525
Density 0.257641 0.001189 0.000679
V Experiments
@@
Table II: OOD recommendation performance comparison on three datasets with seven baselines. Especially, the top two results are emphasized in bold and underlined, denoting the best and second-best results, respectively. Besides, $\dagger$ denotes results are statistically significant where the p-value is less than 0.5.
@@
Metricc R@10 R@20 N@10 N@20 R@50 R@100 N@50 N@100 R@50 R@100 N@50 N@100
FM 5.72 10.74 6.04 7.92 1.21 2.05 0.43 0.57 9.64 13.89 3.13 3.85
NFM 4.05 7.61 4.38 5.60 2.33 3.54 0.66 0.85 8.29 12.76 2.41 3.16
MultiVAE 2.08 4.08 1.72 2.57 2.38 3.68 0.69 0.91 3.65 5.82 1.18 1.54
MacridVAE 2.31 3.92 1.92 2.62 2.19 3.64 0.67 0.90 4.08 6.34 1.35 1.74
MacridVAE+FM 4.63 8.36 5.13 6.43 2.33 3.64 0.66 0.87 4.07 6.26 1.40 1.78
CausPref 5.41 4.77 6.74 9.41 3.44 4.35 0.98 1.21 13.46 18.34 4.14 5.01
COR 7.67 14.43 8.04 10.56 3.68 5.78 1.01 1.35 14.16 19.86 5.00 5.95
Ours 8.20 14.84 8.42 10.86 4.15 5.83 1.13 1.40 14.41 19.90 5.10 6.02
%Improvement 6.91% 2.84% 4.73% 2.84% 12.77% 0.87% 11.88% 3.70% 1.77% 0.20% 2.00% 1.18%
@@
V-A Datasets & Metric
V-A1 Benchmark Datasets.
Following COR [11], we conduct extensive experiments to evaluate the effectiveness of DT3OR on three benchmark datasets, including Meituan  [^1] https://www.biendata.xyz/competition/smp2021_2/, Yelp  [^2] https://www.yelp.com/dataset., and one synthetic dataset.
1) The synthetic dataset contains 1,000 users and 1,000 items. The features of the users/items are sampled from the standard Gaussian $\mathcal{N}(0,1)$. The OOD features are collected by re-sampling user/item features from $\mathcal{N}(1,1)$. The detailed description of the dataset construction is provided in Appendix. LABEL:Constrcution.
2) Meituan includes a public food recommendation dataset that contains a large number of user and item features, e.g., user consumption levels and food prices. For the OOD scenario in Meituan, we define it as a shift in the average consumption levels between weekdays and weekends. Specifically, users exhibit higher or lower consumption levels during weekends compared to weekdays. The user interactions during weekdays are considered in-distribution interactions, while those during weekends are considered OOD interactions.
3) Yelp is a commonly utilized dataset for restaurant recommendations. Within this dataset, the user’s location is taken into account as the fluctuating feature. The user ratings are sorted by timestamps, and the dataset is divided into two parts based on changes in location. The user interactions with shifted locations are considered OOD interactions.
For all datasets, interactions that have ratings of 4 or higher are treated as positive instances. To maintain a level playing field, a sparse partitioning approach is employed, allocating 80%/10%/10% of the data for training, validation, and testing. Comprehensive statistics of the datasets used in this paper are listed in Table I.
V-A2 Evaluation Metrics.

### v2

@@ Abstract
Deep learning has been widely applied in recommender systems, which has recently achieved revolutionary progress. However, most existing learning-based methods assume that the user and item distributions remain unchanged between the training phase and the test phase. However, the distribution of user and item features can naturally shift in real-world scenarios, potentially resulting in a substantial decrease in recommendation performance. This phenomenon can be formulated as an Out-Of-Distribution (OOD) recommendation problem. To address this challenge, we propose a novel Dual Test-Time-Training framework for OOD Recommendation, termed DT3OR. In DT3OR, we incorporate a model adaptation mechanism during the test-time phase to carefully update the recommendation model, allowing the model to adapt specially to the shifting user and item features. To be specific, we propose a self-distillation task and a contrastive task to assist the model learning both the user’s invariant interest preferences and the variant user/item characteristics during the test-time phase, thus facilitating a smooth adaptation to the shifting features. Furthermore, we provide theoretical analysis to support the rationale behind our dual test-time training framework. To the best of our knowledge, this paper is the first work to address OOD recommendation via a test-time-training strategy. We conduct experiments on five datasets with various backbones. Comprehensive experimental results have demonstrated the effectiveness of DT3OR compared to other state-of-the-art baselines.
@@ I Introduction
Recommender systems play a crucial role in alleviating the information overload on social media platforms by providing personalized information filtering. In recent years, a plethora of recommendation algorithms have been proposed, including collaborative filtering [1, 2, 3, 4, 5, 6, 7, 8, 9], graph-based recommendation [10, 11, 12, 13, 14], cross-domain recommendation [15, 16, 17, 18] and etc.
@@
Although promising performance has been achieved, most existing recommendation algorithms assume that the user and item distributions remain unchanged between the training phase and the test phase, adhering to the assumption of independent and identically distributed (I.I.D.) variables. However, these distributions often undergo shifts in many real-world scenarios, violating this assumption. We illustrate the scenario of distribution drift between the training set and the test set in Fig. 1. The shifts primarily manifest in two aspects: user feature shift and item feature shift. User feature shift refers specifically to the evolution of a user’s preferences over time as their individual characteristics change. Specifically, we consider that as individual characteristics evolve over time, there will be corresponding changes in user preferences. For example, as a boy matures, his interests may shift from comic books to sports. Similarly, as he grows older and experiences an increase in income, purchasing a house may become a significant focus during adulthood. In contrast, item feature shift pertains to changes in preference relationships caused by the replacement or updating of item characteristics. For instance, items frequently undergo modifications that alter their features, which may result in outdated user-item interactions and inappropriate recommendations. Consider a restaurant whose signature dishes vary with the seasons: in spring, fresh seasonal vegetables may dominate the menu; in summer, refreshing beverages; in autumn, an assortment of fruits; and in winter, hot pot dishes. Therefore, it is necessary to adapt the recommendation strategy to align with the changing characteristics of the restaurant (item). We demonstrate the user/item feature shift in Fig. 2. These shifts significant impact on recommendation performance [19, 20, 21, 22]. Consequently, it is crucial to consider out-of-distribution (OOD) recommendation to address this issue.
Figure 2: Illustration of user/item feature shift.
The out-of-distribution (OOD) problem has been extensively studied in various fields, including image classification [23] and detection [24]. However, it has received limited attention in the context of recommendation systems. To address this issue, disentangled recommendation approaches [25, 26] aim to learn factorized representations of user preferences, which can enhance the model robustness against distribution shifts. Causal-based methods [19, 20] use causal learning to solve OOD problems. However, these methods often require interventions during the training process, which makes them less practical when only a pre-trained model is available. Another approach to address OOD recommendation is model retraining [27], which involves adapting the model to the OOD environment. However, this approach often requires a substantial computational overhead [19]. In contrast, our study aims to investigate practical approaches that can readily improve the generalization and robustness for a wide range of pre-trained models in different scenarios. Test-time training (TTT), a widely adopted strategy for solving OOD problems in various domains [28], serves as the inspiration for our proposed solution. The core concept of test-time training is to utilize auxiliary unsupervised or self-supervised learning (SSL) tasks to provide insights into the underlying characteristics of a test sample. This information is then used to fine-tune or calibrate the trained model specifically for that test sample. Since each sample has distinct properties, test-time training has significant potential to enhance the generalization capabilities of trained models.
@@
  • In this work, we design test-time training strategies that enable the model to adapt to shifts in feature distributions, addressing the out-of-distribution (OOD) recommendation challenge.
  • In our paper, two self-supervised strategies are introduced to capture users’ invariant preferences, i.e., self-distillation task and contrastive task. To the best of our knowledge, this is the first study to leverage test-time training for addressing the OOD recommendation problem.
  • We provide a theoretical analysis of the effectiveness of the proposed test-time training strategy design.
  • A comprehensive set of comparative experiments on five datasets confirms the effectiveness of our proposed approach relative to existing state-of-the-art methods.
@@ II Problem Definition
III Preliminary
This work proposes test-time training strategies to assist the model in adapting to out-of-distribution scenarios during the test phase. Let $\textbf{E}_{u}$ and $\textbf{E}_{i}$ represent the embeddings of users and items, respectively, obtained from the user embedding extraction network $f(\cdot)$ and the item embedding extraction network $g(\cdot)$ in the latent space. E denotes the fused embeddings of users and items. Additionally, D represents the historical interactions. We use Q and P to denote the original and sharpened distributions of pseudo-labels generated through the clustering operation. R denotes the disjoint clusters, while C represents the user preference clusters. The basic notations are summarized in Tab. I.
IV Methodology
@@
We design a novel framework termed DT3OR for recommendation under an OOD scenario in this section. The framework of DT3OR is shown in Fig. 3. The primary obstacle in optimizing the DT3OR problem lies in constructing a surrogate loss that can guide the learning process, particularly considering the lack of ground-truth interactions for the test data. To be specific, we first delve into the designed self-supervised learning tasks, which include the self-distillation task and the contrastive task. Then, we give the definition of loss function in DT3OR.
IV-A Self-supervised Learning Tasks Design
@@
Our proposed DT3OR is designed to improve model generalizability and robustness to handle OOD recommendations by learning from test data. Ideally, if we had the ground-truth labels for the test data, we could refine the model to minimize the cross-entropy loss on the test data. However, due to the absence of such label information at the test-time, we are motivated to explore suitable surrogate loss functions to steer the model without the presence of labeled data. In the unsupervised scenario, self-supervised learning (SSL) techniques [29, 30, 31] have shown great potential, laying the groundwork for self-supervision during test-time training. The effectiveness of test-time training is heavily dependent on the quality SSL task [28]. However, SSL tasks commonly used in image-related tasks, such as rotation invariance, may not be directly applicable to recommendation data due to its distinct characteristics. Therefore, it is vital to build up a reasonable SSL task based on the characteristics of recommendation scenarios. In recommendation, user preferences often exhibit clustering and homogeneity, where users with similar purchasing behaviors tend to have comparable profiles and are likely to be interested in similar new items. Inspired by this intuition, we designed a dual self-supervised task to assist the model to adapt to the shift distribution in test-time phase, achieving improved recommendation performance on the test set. In following, we design two SSL tasks that are suitable for recommendation tasks. The framework for designed SSL tasks is illustrated in Fig. 4.
Table I: Notation Summary.
Notation Meaning
$\textbf{X}_{u}\in\mathbb{R}^{N_{U}\times d^{\prime}}$ Features of users
$\textbf{X}_{i}\in\mathbb{R}^{N_{I}\times d^{\prime}}$ Features of items
$\textbf{D}\in\mathbb{R}^{N_{D}\times d}$ Historical interactions
$\textbf{E}_{u}\in\mathbb{R}^{N_{D}\times d}$ Embeddings of users
$\textbf{E}_{i}\in\mathbb{R}^{N_{D}\times d}$ Embeddings of items
$\textbf{E}\in\mathbb{R}^{N_{D}\times 2d}$ Fused embeddings of items and users.
$\textbf{Q}\in\mathbb{R}^{N_{D}\times K}$ Distribution of pseudo labels
$\textbf{P}\in\mathbb{R}^{N_{D}\times K}$ Sharpened distribution
$\textbf{R}_{i}\in\mathbb{R}^{N_{D}}$ Clusters of user preference
$\textbf{C}_{i}\in\mathbb{R}^{K}$ Cluster centers
$f(\cdot)$ user embedding extracted network
$g(\cdot)$ item embedding extracted network
IV-A1 Self-distillation Task Design
@@
In this subsection, we propose a test-time self-distillation task to make better leverage of the test data. To be specific, we first extract the embeddings of user/item features and historical interactions by the pretrained user embedding extracted network $f(\cdot)$ and item embedding extracted network $g(\cdot)$ in the training phase:
@@
where $\textbf{X}_{u}$ and $\textbf{X}_{i}$ denote the user features and item features, respectively. D is the historical interactions. $\textbf{E}_{u}$ and $\textbf{E}_{i}$ capture the preferences of users towards items. In a recommendation system, historical interactions and user features are valuable indicators of user preferences [32]. Users’ purchasing behaviors and their interactions with items offer meaningful insights into their underlying preferences.
@@ E $=\gamma(\textbf{E}_{u}\oplus\textbf{E}_{i}),$ (3)
where $\oplus$ denotes the the concatenation operation, $\gamma(\cdot)$ denotes a multi-layer perceptron (MLP). After that, we implement soft K-means algorithm on E. The prediction probability distribution of the fused embedding E could be obtained by:
@@
Here, $\varphi(\cdot)$ represents the soft K-means algorithm, which calculates the assignment probability of each sample to clusters based on its distance from the cluster centers in the latent space. The variable $q$ can be interpreted as the probability of assigning a sample to a specific cluster, indicating a soft assignment. We can consider the matrix $\textbf{Q}=[q]$ as the distribution of assignments for all samples.
@@
IV-A2 Contrastive Task Design
Contrastive learning has surfaced as a significant self-supervised framework, demonstrating promising performance across various tasks [33, 34, 35]. Data augmentation is pivotal in contrastive learning. However, inappropriate data augmentation can lead to semantic drift, thereby degrading the quality of positive samples [36]. Additionally, existing strategies for constructing negative samples often consider all non-positive samples as negative, which can inadvertently introduce false-negative samples [37]. Therefore, we aim to design a contrastive task with more discriminative positive samples and reliable negative samples. The specific process is shown in Fig. 5 The designed contrastive task contains two steps.
Firstly, we construct the positive sample pairs. Based on the clustering distribution $q$ with Eq. 4, we identify the high confidence sample indexes as:
@@ IV-A2 Contrastive Task Design
where $\textbf{C}_{i}$ represents different user preference centers. We consider different high-confidence centers as negative samples, as they correspond to different interest centers. In this way, we could reduce the possibility of false-negative samples.
@@ IV-B Objective function
The first term in Eq. (14) promotes the distribution of user interest centers closer together, enhancing the uniformity of user interest representations in the latent space. We first obtain the probability distribution P of users’ preference centers based on the fused user-item embeddings through a clustering operation. Then, we apply the sharpen function to generate a lower-entropy distribution Q. This distribution reflects the consistency of user preferences in the latent space, indicating that a group of users shares a similar preference for a specific category of items. By minimizing the KL divergence between P and Q, we bring the distributions of user interest centers closer together. The second term captures the correlations among users with similar preferences. Here, we use ${B}$ to represent the batch size. Furthermore, the embeddings’ dimension is set to $d$, and $N$ as the number of high-confidence samples. Consequently, we could obtain time complexity of the self-distillation loss and contrastive loss is $\mathcal{O}(Bd)$ and $\mathcal{O}(N^{2}d)$, respectively. Therefore, the total time complexity of the loss is $\mathcal{O}(Bd+N^{2}d)$. The space complexity of our loss is $\mathcal{O}(B+N^{2})$. The comprehensive learning process of DT3OR is outlined in Algorithm. 1.
V Theoretical analysis
@@
With the inner product of embeddings $g=f(u,i)$ and based on Eq. (15), we could have the following formula:
@@
where $\hat{y}_{i}$ is the prediction of the recommendation model. Then, we calculate the Hessian Matrix of Eq. (17):
@@
Based on Eq. (V), for any ${\bf c}\in\mathbb{R}^{N}$, we could have:
@@
Therefore, we have successfully demonstrated that the Hessian Matrix $\mathcal{\bf M}$ of Eq. (17) is positive semi-definite (PSD). This proof is equivalent to establishing the convexity of $\mathcal{L}(y;{\bf g})$ with respect to $\bf g$.
@@
Remark. By applying Theorem1, we can assert that the objective loss function of the recommendation model possesses the properties of differentiability, convexity, and $\beta$-smoothness. Furthermore, in combination with TheoremV, we can conclude that test-time training could enhance the performance of the recommender system.
@@
Table II: Statistics of the datasets used in our method.
Dataset #User #Item #IID int. #OOD int. Density
Synthetic data 1,000 1,000 145,270 112,371 0.257641
Meituan 2,145 7,189 11,400 6,944 0.001189
Yelp 7,975 74,722 305,128 99,525 0.000679
Amazon-Book 11,000 9,332 75,138 45,326 0.001200
Steam 23,310 5,237 135,961 180,229 0.002600
VI Experiments
@@
  • RQ5: How does the efficiency of DT3OR compare to other recommendation algorithms?
  • RQ6: How stable is DT3OR under distribution shifts of different scales?
Table III: OOD recommendation performance comparison on three datasets with 11 baselines. Especially, the best results are emphasized in bold. Besides, $\dagger$ denotes results are statistically significant where the p-value is less than 0.05.
@@
Metric R@10 R@20 N@10 N@20 R@50 R@100 N@50 N@100 R@50 R@100 N@50 N@100
FM ICDM 2010 5.72 10.74 6.04 7.92 1.21 2.05 0.43 0.57 9.64 13.89 3.13 3.85
NFM SIGIR 2017 4.05 7.61 4.38 5.60 2.33 3.54 0.66 0.85 8.29 12.76 2.41 3.16
MultiVAE WWW 2018 2.08 4.08 1.72 2.57 2.38 3.68 0.69 0.91 3.65 5.82 1.18 1.54
MacridVAE NeurIPS 2019 2.31 3.92 1.92 2.62 2.19 3.64 0.67 0.90 4.08 6.34 1.35 1.74
MacridVAE+FM NeurIPS 2019 4.63 8.36 5.13 6.43 2.33 3.64 0.66 0.87 4.07 6.26 1.40 1.78
CausPref WWW 2022 5.41 4.77 6.74 9.41 3.44 4.35 0.98 1.21 13.46 18.34 4.14 5.01
COR WWW 2022 7.67 14.43 8.04 10.56 3.68 5.78 1.01 1.35 14.16 19.86 5.00 5.95
InvCF WWW 2023 7.64 13.09 7.42 8.45 2.98 4.05 0.74 0.98 10.45 11.86 3.04 4.01
PopGo TOIS 2024 7.08 13.24 7.10 9.52 2.95 4.28 0.85 1.20 12.68 14.62 3.57 4.16
CausalDiffRec Preprint 2024 7.72 14.00 8.09 10.61 3.14 5.29 0.98 1.16 11.98 13.45 3.76 4.38
DR-GNN WWW 2024 7.95 14.21 8.00 10.34 3.72 5.75 1.04 1.35 14.11 19.76 4.87 5.03
Ours 8.20$\dagger$ 14.84$\dagger$ 8.42$\dagger$ 10.86$\dagger$ 4.15$\dagger$ 5.83$\dagger$ 1.13$\dagger$ 1.40$\dagger$ 14.41$\dagger$ 19.90$\dagger$ 5.10$\dagger$ 6.02$\dagger$
%Improvement 3.14% 2.84% 4.08% 2.36% 11.56% 0.87% 8.65% 3.70% 1.77% 0.20% 2.00% 1.18%
@@
VI-A Datasets & Metric
VI-A1 Benchmark Datasets
We conducted extensive experiments to evaluate the effectiveness of DT3OR using five benchmark datasets: Meituan [^1] https://www.biendata.xyz/competition/smp2021_2/, Yelp [^2] https://www.yelp.com/dataset., Steam [^3] https://github.com/kang205/SASRec, Amazon-Book [^4] https://jmcauley.ucsd.edu/data/amazon/, and a synthetic dataset. To comprehensively assess the model, we constructed three distribution shift scenarios: synthetic shift, location shift, and temporal shift. The specific description is detailed as follows.
  • Synthetic Shift. The synthetic dataset contains 1,000 users and 1,000 items. The features of the users/items are sampled from the standard Gaussian $\mathcal{N}(0,1)$. The OOD features are collected by re-sampling user/item features from $\mathcal{N}(1,1)$. The detailed description of the dataset construction is provided in Appendix. IX-B.
  • Location Shift. Yelp is a commonly utilized dataset for restaurant recommendations. Within this dataset, the user’s location is taken into account as the fluctuating feature. The user ratings are sorted by timestamps, and the dataset is divided into two parts based on changes in location. The user interactions with shifted locations are considered OOD interactions.
  • Temporal Shift. Meituan includes a public food recommendation dataset that contains a large number of user and item features, e.g., user consumption levels and food prices. For the OOD scenario in Meituan, we define it as a shift in the average consumption levels between weekdays and weekends. Specifically, users exhibit higher or lower consumption levels during weekends compared to weekdays. The user interactions during weekdays are considered in-distribution interactions, while those during weekends are considered OOD interactions.
  • Temporal Shift. The Amazon Book dataset offers a comprehensive collection of user ratings and reviews for books. It encompasses user IDs, item IDs, ratings, and in-depth reviews, along with extensive metadata including book titles, authors, categories, descriptions, and cover images. With 11,000 users, 9,332 items, and 120,464 interactions (45,326 OOD Int and 75, 138 IID Int). Similar to Meituan dataset, users exhibit higher consumption level during weekends compared with weekdays. Consequently, we classify weekday consumption as in-distribution and weekend consumption as out-of-distribution based on the time of consumption.
  • Temporal Shift. The Steam dataset encompasses textual feedback from users on the Steam platform, featuring contributions from 23,310 users and covering 5,237 items, resulting in a total of 316,190 interactions (180,229 OOD Int and 135,961 IID Int). It includes user reviews, ratings, and diverse metadata related to games, such as titles, genres, and release dates. During weekdays, leisure time is limited, whereas weekends are better suited for relaxation. Similar to the case of Amazon Books, we also separate game time on Steam into two categories: weekday gaming time is considered in-distribution, while weekend gaming time is treated as out-of-distribution.
For all datasets, interactions that have ratings of 4 or higher are treated as positive instances. To maintain a level playing field, a sparse partitioning approach is employed, allocating 80%/10%/10% of the data for training, validation, and testing. Comprehensive statistics of the datasets used in this paper are listed in Tab. II.
VI-A2 Evaluation Metrics